# On Variants of Inverse Cluster Size Problem & General Magnification

### v1

@@ 3.2. Weak General Magnification
Clearly Weak General Magnification implies Weak Cluster Magnification. But the converse is false. Consider the case in Remark 2.4.3. Consider $1\leq j<k\leq n-2$. By proof of Theorem 3 in [krithika2023root], we have $r_{K}(L_{k})=k!$ and $r_{K}(L_{j})=j!$. Thus $s_{K}(L_{k})=\ {n\choose k}$ and $s_{K}(L_{j})=\ {n\choose j}$. Thus $L_{k}/K$ is obtained by nontrivial weak cluster magnification from $L_{j}/K$ but is not in general obtained from non trivial weak general magnification from $L_{j}/K$. For example ${4\choose 1}\nmid\ {4\choose 2}$.
@@
Consider the case in above Remark for k>n/2andj=nkk>n/2\ $\text${and}\ j=n-k. Then $s_{K}(L_{k})=s_{K}(L_{j})$. Also $t_{K}(L_{k})=t_{K}(L_{j})=1$ by Theorem 7.3.6 in [Bhagwat_2025]. Hence we also have that $u_{K}(L_{j})=[L_{j}:K]$ divides $L_{k}:K]=u_{K}(L_{k})$. Hence $L_{k}/K$ is obtained by nontrivial weak general magnification from $L_{j}/K$ but is not obtained by strong general magnification from $L_{j}/K$ since $\tilde{L_{j}}\cap L_{k}=L_{k}\neq L_{j}$ (which contradicts Proposition 3.2 (1)).
@@ 3.3. Strong General Magnification and Unique Chains
    By Theorem 7.1.1 (1) in [Bhagwat_2025], unique intermediate extension in unique descending chain corresponds to subextension fixed by the normaliser subgroup. Now by proof of Proposition 3.3, $N_{G^{\prime}}(H^{\prime})=N_{G}(H)\times N_{S}(T)$. Since $\tilde{L}^{N_{G}(H)}=N_{1}$ and $\tilde{J}^{N_{S}(T)}=N^{\prime}_{1}$. Thus $\tilde{M}^{N_{G^{\prime}}(H^{\prime})}=N_{1}N^{\prime}_{1}$.
@@
    By Theorem 7.2.1 (3) in [Bhagwat_2025], unique intermediate extension in unique ascending chain corresponds to subextension fixed by the normal closure subgroup. Now by proof of Proposition 3.3, $H^{\prime G^{\prime}}=H^{G}\times T^{S}$. Since $\tilde{L}^{H^{G}}=F_{1}$ and $\tilde{J}^{T^{S}}=F^{\prime}_{1}$. Thus $\tilde{M}^{H^{\prime G^{\prime}}}=F_{1}F^{\prime}_{1}$.
@@
    We have $N_{i}\neq L$ but $N^{\prime}_{i}=J$ and we have $N^{\prime}_{i}\neq K$ but $N_{i}=K$. Now clearly $J=N^{\prime}_{i}\neq K$. Thus the magnification is nontrivial. Also $i>0$ as $N_{i}\neq L$. Since $N^{\prime}_{i}=J$ for $i>0$, we have that unique descending chain of $J/K$ is singleton $J=N^{\prime}_{0}$. Hence by Theorem 7.1.1 in [Bhagwat_2025], we get $r_{K}(J)=1$. Now since $N_{i}=K$, so the unique descending chain of $L/K$ terminates at $K$. Now $N_{i}=F_{j}$ and $N^{\prime}_{i}=F^{\prime}_{j}$. Thus we have $F_{j}\neq L$ but $F^{\prime}_{j}=J$ and we have $F^{\prime}_{j}\neq K$ but $F_{j}=K$. Now $j>0$ as $F^{\prime}_{j}\neq K$. Since $F_{j}=K$ for $j>0$, we have that unique ascending chain of $L/K$ is singleton $K=F_{0}$. Hence by Theorem 7.2.1 in [Bhagwat_2025], we get $t_{K}(L)=1$. Now since $F^{\prime}_{j}=J$, so the unique ascending chain of $J/K$ terminates at $J$.
@@ 4.1. General Primitive Extensions
We claim that $LJ/\mathbb{Q}$ is primitive. We have that Galois closure of $LJ/\mathbb{Q}$ is $\tilde{L}\tilde{J}$. We also have that $Gal(\tilde{L}/L)$ and $Gal(\tilde{J}/J)$ are isomorphic to $\mathfrak{A}_{n-1}$. Thus $Gal(\tilde{L}\tilde{J}/\mathbb{Q})\cong\mathfrak{A}_{n}\times\mathfrak{A}_{n}$ and $Gal(\tilde{L}\tilde{J}/LJ)\cong\mathfrak{A}_{n-1}\times\mathfrak{A}_{n-1}$. Since $\mathfrak{A}_{n}$ is simple non-abelian for $n\geq 5$. Thus the only proper nontrivial normal subgroups of $\mathfrak{A}_{n}\times\mathfrak{A}_{n}$ are $1\times\mathfrak{A}_{n}$ and $\mathfrak{A}_{n}\times 1$ both of which do not contain $\mathfrak{A}_{n-1}\times\mathfrak{A}_{n-1}$. Hence by Proposition 2.3 we are done.
@@
Observe that if $L/K$ is Galois then by Proposition 3.2 (5) and Corollary 4.1.8 in [Bhagwat_2025], $L/K$ is primitive if and only if $L/K$ is general primitive.
@@
Consider $L/K$. Let $G=Gal(\tilde{L}/K)$. If $G$ has less than two proper nontrivial normal subgroups then $L/K$ is general primitive. In particular if $G$ is simple then $L/K$ is general primitive.
@@ 4.2. General Primitive Inverse Cluster Size Problem
Note that Theorem 2.6 was the Primitive variant of the problem for the cases $(K,n,r)$ where $r<n$ which was solved by the approach of Theorem 3.1.1 in [Bhagwat_2025] in the light of Theorem 2.4. But the General variant of the problem cannot be easily tackled with the same approach because of the conclusions of Theorem 3.9. So in this article we take a case by case approach for the General variant and establish some interesting cases.
@@
By results in [volklein1996groups] on hilbertian fields, we have ${\mathfrak{S}}_{n}$ to be realisable as a Galois group over $K$. Thus by the final proposition in [perlisroots], there exist polynomial $f$ over $K$ of degree $n$ with Galois group ${\mathfrak{S}}_{n}$. Suppose the roots of $f$ are $\alpha_{i}\in\bar{K}$ for $1\leq i\leq n$. For $1\leq k\leq n-2$, let $L_{k}=K(\alpha_{1},\dots,\alpha_{k})$. From Theorem 3 in [krithika2023root], we have $L_{k}:K]=\ ^{n}P_{k}$ and $r_{K}(L_{k})=k!$.
Now Galois closure for each $L_{k}/K$ for $1\leq k\leq n-2$ is $L_{n-1}$ with $Gal(L_{n-1}/K)$ as $\mathfrak{S}_{n}$. For $n\neq 4$, there is only one proper nontrivial normal subgroup of $\mathfrak{S}_{n}$. Hence we are done by Proposition 4.2. For $n=4$, there are two proper nontrivial normal subgroups of $\mathfrak{S}_{4}$ but one is contained in the other hence $L_{k}/K$ is general primitive by Proposition 3.2 (2).∎
@@
Now the case $(K,4,2)$. The dihedral group $D_{4}$ of order $8$ is solvable hence is realizable over any number field by Lemma 6.6 (2) in [jaiswal2025rootcapacityintersectionindicium]. Thus by the final proposition in [perlisroots], there exists $L/K$ with degree $4$ with Galois group of Galois closure as $D_{4}$. It is easy to see that $r_{K}(L)=2$. Now since intersection of any two nontrivial normal subgroups of $D_{4}$ is nontrivial (as the center of $D_{4}$, a subgroup of order $2$, is contained there), $L/K$ is general primitive by Proposition 3.2 (2).∎
@@
Since $r\geq 3$, we can assume $p\geq 5$. By Theorem 1.4 in [zywina2023modular], $PSL_{2}(\mathbb{F}_{p})$ is realizable as a Galois group over $\mathbb{Q}$ for all primes $p\geq 5$. For any prime $p\geq 5$, let $G=PSL_{2}(\mathbb{F}_{p})$. Now |G|=(p1)p(p+1)/2|G|=(p-1)p(p+1)/{2}.
@@
Consider the normalizer of $H$ in $B_{2}(\mathbb{F}_{p})$ which is $N_{B_{2}(\mathbb{F}_{p})}(H)\cong\{0\}\rtimes\mathbb{F}_{p}^{\times}$ since k=(p1)/r>2k=(p-1)/r>2. So $N_{G}(H):H]=r$ and $[G:H]=pr$. Since $G$ is solvable hence is realizable over any number field by Lemma 6.6 (2) in [jaiswal2025rootcapacityintersectionindicium]. Thus we have existence of $L/K$ (precisely the fixed field of $H$) with given degree and cluster size. We also have $\cap_{g\in G}\ gHg^{-1}=\{1\}$ as $k>2$. Thus the Galois group of Galois closure of $L/K$ is $G$ itself.
@@
We also have that elements of $B$ centralise elements of $A$ and that $p\nmid|B|$ and $A\cap B=1$. Thus one can show that B=<(0,1)><I>B=\ <(0,-1)>\ $\cong\$<-I>. Thus |C|=(p1)/2|C|=(p-1)/2. We know that unique index-$2$ subgroup of $\mathbb{F}_{p}^{\times}$ is $(\mathbb{F}_{p}^{\times})^{2}$. Thus $C=(\mathbb{F}_{p}^{\times})^{2}$. Now consider the case $p\equiv 1\pmod{4}$. We have that $1\in C$. Thus $(0,-1)\in A$ which contradicts $A\cap B=1$.

### v2

@@ 3.2. Weak General Magnification
Clearly Weak General Magnification implies Weak Cluster Magnification. But the converse is false. Consider the case in Remark 2.4.3. Consider $1\leq j<k\leq n-2$. By proof of Theorem 3 in [9], we have $r_{K}(L_{k})=k!$ and $r_{K}(L_{j})=j!$. Thus $s_{K}(L_{k})=\ {n\choose k}$ and $s_{K}(L_{j})=\ {n\choose j}$. Thus $L_{k}/K$ is obtained by nontrivial weak cluster magnification from $L_{j}/K$ but is not in general obtained from non trivial weak general magnification from $L_{j}/K$. For example ${4\choose 1}\nmid\ {4\choose 2}$.
@@
Consider the case in above Remark for k>n/2andj=nkk>n/2\ $\text${and}\ j=n-k. Then $s_{K}(L_{k})=s_{K}(L_{j})$. Also $t_{K}(L_{k})=t_{K}(L_{j})=1$ by Theorem 7.3.6 in [1]. Hence we also have that $u_{K}(L_{j})=[L_{j}:K]$ divides $L_{k}:K]=u_{K}(L_{k})$. Hence $L_{k}/K$ is obtained by nontrivial weak general magnification from $L_{j}/K$ but is not obtained by strong general magnification from $L_{j}/K$ since $\tilde{L_{j}}\cap L_{k}=L_{k}\neq L_{j}$ (which contradicts Proposition 3.2 (1)).
@@ 3.3. Strong General Magnification and Unique Chains
    By Theorem 7.1.1 (1) in [1], unique intermediate extension in unique descending chain corresponds to subextension fixed by the normalizer subgroup. Now by proof of Proposition 3.3, $N_{G^{\prime}}(H^{\prime})=N_{G}(H)\times N_{S}(T)$. Since $\tilde{L}^{N_{G}(H)}=N_{1}$ and $\tilde{J}^{N_{S}(T)}=N^{\prime}_{1}$. Thus $\tilde{M}^{N_{G^{\prime}}(H^{\prime})}=N_{1}N^{\prime}_{1}$.
@@
    By Theorem 7.2.1 (3) in [1], unique intermediate extension in unique ascending chain corresponds to subextension fixed by the normal closure subgroup. Now by proof of Proposition 3.3, $H^{\prime G^{\prime}}=H^{G}\times T^{S}$. Since $\tilde{L}^{H^{G}}=F_{1}$ and $\tilde{J}^{T^{S}}=F^{\prime}_{1}$. Thus $\tilde{M}^{H^{\prime G^{\prime}}}=F_{1}F^{\prime}_{1}$.
@@
    We have $N_{i}\neq L$ but $N^{\prime}_{i}=J$ and we have $N^{\prime}_{i}\neq K$ but $N_{i}=K$. Now clearly $J=N^{\prime}_{i}\neq K$. Thus the magnification is nontrivial. Also $i>0$ as $N_{i}\neq L$. Since $N^{\prime}_{i}=J$ for $i>0$, we have that unique descending chain of $J/K$ is singleton $J=N^{\prime}_{0}$. Hence by Theorem 7.1.1 in [1], we get $r_{K}(J)=1$. Now since $N_{i}=K$, so the unique descending chain of $L/K$ terminates at $K$. Now $N_{i}=F_{j}$ and $N^{\prime}_{i}=F^{\prime}_{j}$. Thus we have $F_{j}\neq L$ but $F^{\prime}_{j}=J$ and we have $F^{\prime}_{j}\neq K$ but $F_{j}=K$. Now $j>0$ as $F^{\prime}_{j}\neq K$. Since $F_{j}=K$ for $j>0$, we have that unique ascending chain of $L/K$ is singleton $K=F_{0}$. Hence by Theorem 7.2.1 in [1], we get $t_{K}(L)=1$. Now since $F^{\prime}_{j}=J$, so the unique ascending chain of $J/K$ terminates at $J$.
@@ 4.1. General Primitive Extensions
We claim that $LJ/\mathbb{Q}$ is primitive. We have that Galois closure of $LJ/\mathbb{Q}$ is $\tilde{L}\tilde{J}$. We also have that ${{\rm Gal}}(\tilde{L}/L)$ and ${{\rm Gal}}(\tilde{J}/J)$ are isomorphic to $\mathfrak{A}_{n-1}$. Thus ${{\rm Gal}}(\tilde{L}\tilde{J}/\mathbb{Q})\cong\mathfrak{A}_{n}\times\mathfrak{A}_{n}$ and ${{\rm Gal}}(\tilde{L}\tilde{J}/LJ)\cong\mathfrak{A}_{n-1}\times\mathfrak{A}_{n-1}$. Since $\mathfrak{A}_{n}$ is simple non-abelian for $n\geq 5$. Thus the only proper nontrivial normal subgroups of $\mathfrak{A}_{n}\times\mathfrak{A}_{n}$ are $1\times\mathfrak{A}_{n}$ and $\mathfrak{A}_{n}\times 1$ both of which do not contain $\mathfrak{A}_{n-1}\times\mathfrak{A}_{n-1}$. Hence by Proposition 2.3 we are done.
@@
Observe that if $L/K$ is Galois then by Proposition 3.2 (5) and Corollary 4.1.8 in [1], $L/K$ is primitive if and only if $L/K$ is general primitive.
@@
Consider $L/K$. Let $G={{\rm Gal}}(\tilde{L}/K)$. If $G$ has less than two proper nontrivial normal subgroups then $L/K$ is general primitive. In particular if $G$ is simple then $L/K$ is general primitive.
@@ 4.2. General Primitive Inverse Cluster Size Problem
Note that Theorem 2.6 was the Primitive variant of the problem for the cases $(K,n,r)$ where $r<n$ which was solved by the approach of Theorem 3.1.1 in [1] in the light of Theorem 2.4. But the General variant of the problem cannot be easily tackled with the same approach because of the conclusions of Theorem 3.9. So in this article we take a case by case approach for the General variant and establish some interesting cases. We begin with some lemmas.
Lemma 4.3 (Final Proposition , [11]).
Let $G$ be a transitive subgroup of ${\mathfrak{S}}_{n}$ for some $n$. If there exists a finite Galois extension of a field $K$ with Galois group isomorphic to $G$, then there exists an irreducible polynomial $f$ over $K$ of degree $n$ and a labelling of the roots of $f$ so that the Galois group of $f$, viewed as a group permuting roots of $f$, is precisely $G$.
Lemma 4.4 (Lemma 6.6 (2), [5]).
Let $K$ be a number field. Given $L/K$, for any solvable group $G$, we have a Galois extension $L^{\prime}/K$ such that ${{\rm Gal}}(L^{\prime}/K)=G$ and $L^{\prime}\cap L=K$.
@@
By results in [13] on hilbertian fields, we have ${\mathfrak{S}}_{n}$ to be realisable as a Galois group over $K$. Thus by Lemma 4.3, there exist polynomial $f$ over $K$ of degree $n$ with Galois group ${\mathfrak{S}}_{n}$. Suppose the roots of $f$ are $\alpha_{i}\in\bar{K}$ for $1\leq i\leq n$. For $1\leq k\leq n-2$, let $L_{k}=K(\alpha_{1},\dots,\alpha_{k})$. From Theorem 3 in [9], we have $L_{k}:K]=\ ^{n}P_{k}$ and $r_{K}(L_{k})=k!$.
Now Galois closure for each $L_{k}/K$ for $1\leq k\leq n-2$ is $L_{n-1}$ with ${{\rm Gal}}(L_{n-1}/K)$ as $\mathfrak{S}_{n}$. For $n\neq 4$, there is only one proper nontrivial normal subgroup of $\mathfrak{S}_{n}$. Hence we are done by Proposition 4.2. For $n=4$, there are two proper nontrivial normal subgroups of $\mathfrak{S}_{4}$ but one is contained in the other hence $L_{k}/K$ is general primitive by Proposition 3.2 (2).∎
@@
Now the case $(K,4,2)$. The dihedral group $D_{4}$ of order $8$ is solvable hence is realizable over any number field by Lemma 4.4. Thus by the Lemma 4.3, there exists $L/K$ with degree $4$ with Galois group of Galois closure as $D_{4}$. It is easy to see that $r_{K}(L)=2$. Now since intersection of any two nontrivial normal subgroups of $D_{4}$ is nontrivial (as the center of $D_{4}$, a subgroup of order $2$, is contained there), $L/K$ is general primitive by Proposition 3.2 (2).∎
@@
Since $r\geq 3$, we can assume $p\geq 5$. By Theorem 1.4 in [14], $PSL_{2}(\mathbb{F}_{p})$ is realizable as a Galois group over $\mathbb{Q}$ for all primes $p\geq 5$. For any prime $p\geq 5$, let $G=PSL_{2}(\mathbb{F}_{p})$. Now |G|=(p1)p(p+1)/2|G|=(p-1)p(p+1)/{2}.
@@
Consider the normalizer of $H$ in $B_{2}(\mathbb{F}_{p})$ which is $N_{B_{2}(\mathbb{F}_{p})}(H)\cong\{0\}\rtimes\mathbb{F}_{p}^{\times}$ since k=(p1)/r>2k=(p-1)/r>2. So $N_{G}(H):H]=r$ and $[G:H]=pr$. Since $G$ is solvable hence is realizable over any number field by Lemma 4.4. Thus we have existence of $L/K$ (precisely the fixed field of $H$) with given degree and cluster size. We also have $\cap_{g\in G}\ gHg^{-1}=\{1\}$ as $k>2$. Thus the Galois group of Galois closure of $L/K$ is $G$ itself.
@@
We also have that elements of $B$ centralize elements of $A$ and that $p\nmid|B|$ and $A\cap B=1$. Thus one can show that B=<(0,1)><I>B=\ <(0,-1)>\ $\cong\$<-I>. Thus |C|=(p1)/2|C|=(p-1)/2. We know that unique index-$2$ subgroup of $\mathbb{F}_{p}^{\times}$ is $(\mathbb{F}_{p}^{\times})^{2}$. Thus $C=(\mathbb{F}_{p}^{\times})^{2}$. Now consider the case $p\equiv 1\pmod{4}$. We have that $1\in C$. Thus $(0,-1)\in A$ which contradicts $A\cap B=1$.